OpenStudy (anonymous):

Hi! I need help with this derivative problem: If a ball is thrown into the air with a velocity of 40 ft/s, its height after t seconds is given by y=10t-16t^2. Find the velocity when t=2

OpenStudy (anonymous):

dy/dx=10-32t

OpenStudy (anonymous):

dy/dx is velocity.as y is the displacement and its derivative i.e. rate of chnge of displacement will be velocity

OpenStudy (anonymous):

sorry it will be dy/dt=10-32t

OpenStudy (anonymous):

when t=2, v=10-32*2 =10-64 =-54ft/s